# ViewSRD: 3D Visual Grounding via Structured Multi-View Decomposition

### v1

@@ Abstract
3D visual grounding aims to identify and localize objects in a 3D space based on textual descriptions. However, existing methods struggle with disentangling targets from anchors in complex multi-anchor queries and resolving inconsistencies in spatial descriptions caused by perspective variations. To tackle these challenges, we propose ViewSRD, a framework that formulates 3D visual grounding as a structured multi-view decomposition process. First, the Simple Relation Decoupling (SRD) module restructures complex multi-anchor queries into a set of targeted single-anchor statements, generating a structured set of perspective-aware descriptions that clarify positional relationships. These decomposed representations serve as the foundation for the Multi-view Textual-Scene Interaction (Multi-TSI) module, which integrates textual and scene features across multiple viewpoints using shared, Cross-modal Consistent View Tokens (CCVTs) to preserve spatial correlations. Finally, a Textual-Scene Reasoning module synthesizes multi-view predictions into a unified and robust 3D visual grounding. Experiments on 3D visual grounding datasets show that ViewSRD significantly outperforms state-of-the-art methods, particularly in complex queries requiring precise spatial differentiation. Code is available at https://github.com/xxx.
00footnotetext: Accepted by ICCV 2025. Current version is not the final version.11footnotetext: The first two authors contributed equally.22footnotetext: Corresponding authors:xuemx@scut.edu.cn, huaidongz@scut.edu.cn.
@@ 1 Introduction
Figure 1: (a) Previous 3DVG methods struggle with ambiguities from complex multi-anchor queries and perspective shifts. (b) ViewSRD addresses this by using the Simple Relation Decoupling (SRD) module to simplify queries and the cross-modal consistent view tokens (CCVTs) to capture viewpoint variations in both scene and textual modal, boosting cross-modal feature interaction and enhancing grounding accuracy.
@@
To tackle these challenges, we propose ViewSRD, a framework that formulates 3D visual grounding as a structured multi-view decomposition process. By leveraging the Simple Relation Decoupling (SRD) module, ViewSRD effectively disentangles target-anchor relationships in the complex multi-anchor queries, while the Multi-view Textual-Scene Interaction (Multi-TSI) module integrates multi-view information to enhance grounding accuracy. As illustrated in Fig. 1(b), ViewSRD first applies the SRD module to decompose complex multi-anchor queries into a set of simpler single-anchor queries, isolating interactions between the target and its anchors. This structured decomposition allows the model to more effectively learn positional relationships from textual descriptions. The Multi-TSI module then fuses textual and scene features across multiple viewpoints using Cross-modal Consistent View Tokens (CCVTs), which explicitly encode viewpoint information as learnable cue for both textual and scene module. This mechanism ensures that the model accurately captures spatial interactions, even under perspective shifts. Finally, the Textual-Scene Reasoning module aggregates these multi-view features to accurately predict the final 3D VG results. Extensive experiments have validated the efficacy of our proposed ViewSRD across different 3DVG benchmarks, demonstrating its superior performance across diverse scenarios.
In summary, our contributions are fourfold:
@@
  • We conduct extensive evaluations on multiple 3D visual grounding datasets, where ViewSRD achieves state-of-the-art performance, surpassing existing methods by a significant margin.
@@ 3 ViewSRD
Both these modules enhance the robustness and accuracy of ViewSRD in multi-view 3DVG.
@@ 3.2 Textual Aggregation
where $\alpha$ is uniformly sampled from $\{0,0.1,0.3,0.5\}$ during training and fixed at 0.5 during validation. This adaptive fusion strategy ensures smooth feature integration, enhancing the model’s robustness in language-conditioned 3D grounding.
@@ 3.5 Overall Loss Functions
Following prior research [35, 18], three distinct loss functions are applied on ViewSRD. These include a referential loss $\mathcal{L}_{Ref}$ derived from grounding predictions, an object-level loss $\mathcal{L}_{Object}$ capturing object shape and center and a sentence-level loss $\mathcal{L}_{Sent}$ designed to identify the target and anchor phrases within the $F_{\text{agg}}$. Similarly to [2], we extend the referential loss $\mathcal{L}_{Ref}$ to localize both the target and the anchors, which we term as parallel referential loss $\mathcal{L}^{P}_{\text{ref}}$, where both the target and anchors are localized simultaneously. For the details of these three losses, please refer to the supplementary materials. The total loss function is defined as:
$\mathcal{L}=\mathcal{L}_{Object}+\mathcal{L}^{P}_{Ref}+\mathcal{L}_{Sent}.$ (7)
@@
4 Experiments
4.1 Experiment Settings
Datasets. Nr3D [1] consists of 45,503 human utterances annotating 707 indoor scenes from the ScanNet [10] database, covering 76 fine-grained object categories with multiple same-category distractors. Sr3D [1] comprises 83,572 templated sentences structured in a “target-space relation-anchor object” format, featuring a simpler setup with a similar distractor configuration as Nr3D. ScanRefer [6] contains 51,583 human-written descriptions for 11,046 objects in 800 ScanNet [10] scenes, incorporating spatial relations and object attributes to facilitate 3DVG.
Evaluation Metrics. For Nr3D and Sr3D datasets [1], accuracy performance is evaluated by measuring the percentage of successful matches boxes, as commonly done in previous works [18, 35]. For the ScanRefer dataset [6], accuracy is evaluated as the percentage of correctly predicted bounding boxes whose IoU (Intersection over Union) with the ground truth exceeds 0.25 (Acc@0.25) or 0.5 (Acc@0.5), following standard evaluation protocols [41].
Implementation details. We use Pytorch for our experiment on a GeForce RTX 4090 GPU. We use AdamW [29, 46, 22, 42] as the optimizer with a learning rate of 0.0005 in our experiment. We set the number of views N to 4. We use the DeepSeek-R1 [27] as the LLM in the SRD module, as it offers a strong balance between performance and reproducibility, and can be efficiently deployed on a single RTX 4090 GPU.
4.2 3D Visual Grounding Results
In this section, we conduct a comparative experiment between the ViewSRD and state-of-the-art methods to assess the advantages of our proposed approach for 3DVG.
We first analyze some complex-query cases from Nr3D [1] in Fig.4, where we visualized the grounding truth boxes, MVT [18] predictions, CoT3DRef [2] predictions, ViewSRD predictions, query sentences, and simple sentences decoupled by the SRD Module along with their corresponding target-anchor pairs in the 3D scene. We observed that MVT and CoT3DRef fail to ensure the correct object alignment in cases where the query sentence involves multiple anchors, such as in the case with “bed”, “table”, and “chair”. In contrast, ViewSRD successfully grounds the correct target by understanding the spatial relations between the target-anchor pairs and leveraging the learning from simple sentences. Furthermore, Under the enhanced strategy of viewpoint rotation, CoT3DRef fails to ensure the correct viewpoint alignment as described in the sentence and grounds to the wrong object. On the other hand, ViewSRD successfully predicts the target by leveraging the spatial relationship of the target-anchor pairs that are invariant to viewpoint changes, such as “The trash cans are under the sink”, leading to more stable grounding compared to target-anchor pairs susceptible to viewpoint variations.
The Quantitative results on Nr3D [1] are shown in Table 1, our ViewSRD outperforms all previous methods. For example, ViewSRD surpasses the best competitor, CoT3DRef [2], with a 5.2% improvement in overall accuracy performance under identical training conditions. In the comparison of experimental results under view dependency conditions, our method outperforms CoT3DRef [2] 6.7%. These experimental results show that CCVTs guide the model to dynamically adjust both the textual and visual spaces. Through the decoupled sentences ViewSRD’s ability to understand whether the multiple spatial relationships mentioned in the sentence exhibit viewpoint dependency.
In Table 1, we further present experimental results on Sr3D [1] to verify the generalizability of our approach. It can be seen that our method performs exceptionally well across all evaluation metrics, particularly achieving the highest accuracy of 76.2% in the View Independent scenario, demonstrating the model’s strong generalization ability in handling tasks that do not rely on viewpoint. Additionally, our method achieved +2.8% and +2.7% in the View Independence and Hard scenarios, respectively, further validating its robustness in complex and challenging tasks.
Figure 5: t-sne Visualization of CCVTs.
To further understand how CCVTs encode perspective information, we visualize the learned CCVT features using t-SNE, as shown in Figure 5. Tokens corresponding to different viewpoints form distinct clusters, while those from the same viewpoint appear closely grouped. This clearly demonstrates that CCVTs capture and preserve view-specific semantics, which are crucial for accurate alignment in 3D visual grounding.
4.3 Analysis of Anchors on the SR3D/NR3D.
@@ 4.1 Experiment Settings
Table 3: Performance (%) of SRD module improves MVT [18] and BUTD-DETR [21] on ScanRefer [6] dataset.
@@ 4.5 Ablation Study
Analysis of ViewSRD Components. To assess the contribution of each component within ViewSRD, we conducted detailed ablation studies on the Nr3D dataset [1]. Starting from the full model, we systematically removed key modules one at a time to evaluate their individual impact. The results, presented in Table 4, demonstrate that each component plays a crucial role in enhancing model performance across different scenarios. Notably, the removal of the CCVTs leads to the most significant performance degradation. This is primarily because, without the view token, the model lacks explicit viewpoint information, impairing its ability to distinguish between different perspectives. Similarly, removing either the textual module or the scene module results in a noticeable decline in performance, underscoring the necessity of cross-modal interaction. When view-alignment weighting is disabled (w/o Weight), performance drops by 0.9%, showing that dynamic alignment of view features is critical for performance under view-dependent conditions. Removing the SRD module leads to performance degradation, confirming the benefit of multi-anchor query decoupling. We also compare it with an LLM-based augmentation method from Multi3DRefer and find that SRD achieves greater gains, highlighting the advantage of structured sub-query decomposition over generic augmentation.
Analysis of Multi-View Modeling. We further investigate the impact of varying the number of views on 3DVG performance during both training and testing on the Nr3D dataset. As shown in Table 5, increasing the number of views during testing consistently enhances performance when the model is trained with four views. This trend underscores the effectiveness of multi-view aggregation in improving grounding accuracy by leveraging complementary spatial information from multiple perspectives.
Additionally, performance variations emerge when the number of views remains consistent between training and testing. Training and testing with a single view yield an accuracy of 64.4%, which increases to 67.7% with two views. However, further increasing the number of views to eight results in only a marginal gain, reaching 68.4%. These results indicate that while incorporating additional viewpoints benefits performance, the improvements eventually plateau, suggesting a point of diminishing returns. Notably, using four views provides a well-balanced multi-view representation, effectively capturing diverse spatial information while minimizing redundancy and maintaining fast computation. This suggests that equal attention distribution across many views can dilute focus from crucial perspectives. Future work will explore dynamic view selection strategies.

### v2

@@ Abstract
3D visual grounding aims to identify and localize objects in a 3D space based on textual descriptions. However, existing methods struggle with disentangling targets from anchors in complex multi-anchor queries and resolving inconsistencies in spatial descriptions caused by perspective variations. To tackle these challenges, we propose ViewSRD, a framework that formulates 3D visual grounding as a structured multi-view decomposition process. First, the Simple Relation Decoupling (SRD) module restructures complex multi-anchor queries into a set of targeted single-anchor statements, generating a structured set of perspective-aware descriptions that clarify positional relationships. These decomposed representations serve as the foundation for the Multi-view Textual-Scene Interaction (Multi-TSI) module, which integrates textual and scene features across multiple viewpoints using shared, Cross-modal Consistent View Tokens (CCVTs) to preserve spatial correlations. Finally, a Textual-Scene Reasoning module synthesizes multi-view predictions into a unified and robust 3D visual grounding. Experiments on 3D visual grounding datasets show that ViewSRD significantly outperforms state-of-the-art methods, particularly in complex queries requiring precise spatial differentiation. Code is available at https://github.com/visualjason/ViewSRD.
00footnotetext: Accepted by ICCV 2025.11footnotetext: The first two authors contributed equally.22footnotetext: Corresponding authors:xuemx@scut.edu.cn, harxis@outlook.com.
@@ 1 Introduction
Figure 1: (a) Previous 3DVG methods struggle with ambiguities from complex multi-anchor queries and perspective shifts. (b) ViewSRD addresses this by using the SRD module to simplify queries and the CCVTs to capture viewpoint variations in both scene and textual modal, boosting cross-modal feature interaction and enhancing grounding accuracy.
@@
To tackle these challenges, we propose ViewSRD, a framework that formulates 3D visual grounding as a structured multi-view decomposition process. By leveraging the Simple Relation Decoupling (SRD) module, ViewSRD effectively disentangles target-anchor relationships in the complex multi-anchor queries, while the Multi-view Textual-Scene Interaction (Multi-TSI) module integrates multi-view information to enhance grounding accuracy. As illustrated in Fig. 1(b), ViewSRD first applies the SRD module to decompose complex multi-anchor queries into a set of simpler single-anchor queries, isolating interactions between the target and its anchors. This structured decomposition allows the model to more effectively learn positional relationships from textual descriptions. The Multi-TSI module then fuses textual and scene features across multiple viewpoints using Cross-modal Consistent View Tokens (CCVTs), which explicitly encode viewpoint information as learnable cue for both textual and scene module. This mechanism ensures that the model accurately captures spatial interactions, even under perspective shifts. Finally, the Textual-Scene Reasoning module aggregates these multi-view features to accurately predict the final 3D VG results. Extensive experiments have validated the efficacy of our proposed ViewSRD across different 3DVG benchmarks, demonstrating its superior performance across diverse scenarios. In summary, our contributions are fourfold:
@@
  • We conduct extensive evaluations on 3D visual grounding datasets, where ViewSRD achieves state-of-the-art performance, yielding superior performance over prior work.
@@ 3.2 Textual Aggregation
[…]
@@ 3.5 Overall Loss Functions
Following prior research [36, 18], three distinct loss functions are applied on ViewSRD. These include a referential loss $\mathcal{L}_{Ref}$ derived from grounding predictions, an object-level loss $\mathcal{L}_{Object}$ capturing object shape and center and a sentence-level loss $\mathcal{L}_{Sent}$ designed to identify the target and anchor phrases within the $F_{\text{agg}}$. Similarly to [2], we extend the referential loss $\mathcal{L}_{Ref}$ to localize both the target and the anchors, which we term as parallel referential loss $\mathcal{L}^{P}_{\text{ref}}$, where both the target and anchors are localized simultaneously. For details of these losses, please refer to supplementary materials. The total loss function is defined as:
$\mathcal{L}=\lambda_{Obj}\mathcal{L}_{Object}+\lambda_{Ref}\mathcal{L}^{P}_{Ref}+\lambda_{Sent}\mathcal{L}_{Sent}.$ (7)
4 Experiments
4.1 Experiment Settings
Datasets. Nr3D [1] contains 45,503 human utterances referencing 707 indoor scenes from ScanNet [10], covering 76 object categories with multiple same-class distractors. Sr3D [1] includes 83,572 template-based sentences in a “target–spatial relation–anchor” format, offering a simpler setup with similar distractors. ScanRefer [6] provides 51,583 free-form descriptions for 11,046 objects across 800 ScanNet scenes, incorporating spatial and attribute-level references to support 3DVG.
Evaluation Metrics. For Nr3D and Sr3D, grounding accuracy is measured by the percentage of correctly matched boxes [18, 36]. For ScanRefer, we report Acc@0.25 and Acc@0.5, i.e., the percentage of predicted boxes with IoU exceeding 0.25 or 0.5, respectively [42].
Implementation Details. All experiments are implemented in PyTorch and run on a single RTX 4090 GPU. We use AdamW [29, 47, 22, 43] with a learning rate of 0.0005. The number of input views is set to $N=4$. We set $\lambda_{Ref},\lambda_{Obj},\lambda_{Sent}=1.0,0.5,0.5$. For the SRD module, we adopt DeepSeek-R1 [27], which balances performance and reproducibility, and can runs on a RTX 4090.
4.2 3D Visual Grounding Results
@@ 4.1 Experiment Settings
We compare ViewSRD with recent state-of-the-art approaches to evaluate its effectiveness on 3DVG. Fig. 4 illustrates complex query cases from Nr3D [1], including ground truth boxes, predictions from MVT [18], CoT3DRef [2], and ViewSRD, along with the original queries and the simplified sentences produced by the SRD module. In multi-anchor scenarios (e.g., involving “bed”, “table”, and “chair”), MVT and CoT3DRef often misalign predictions due to challenges in spatial reasoning. In contrast, ViewSRD correctly grounds targets by decomposing complex queries and leveraging robust spatial relationships between target-anchor pairs. Moreover, under viewpoint shifts, CoT3DRef struggles to maintain alignment, whereas ViewSRD reliably grounds targets by capturing spatial relations invariant to viewpoint changes (e.g., “The trash cans are under the sink”).
Quantitative results on Nr3D (Table 1) show that ViewSRD achieves a 5.2% accuracy gain over the best prior method, CoT3DRef, under identical settings. Under view-dependent evaluation, it further outperforms CoT3DRef by 6.7%, demonstrating the effectiveness of CCVTs in aligning textual and visual spaces and modeling viewpoint-sensitive relations through query decomposition. To assess generalization, we also evaluate on Sr3D [1] (Table 1). ViewSRD achieves the highest accuracy of 76.2% in the View-Independent setting, with additional gains of +2.8% and +2.7% in the View-Independence and Hard scenarios, respectively. These results confirm the robustness and generalizability of our approach across diverse scenario.
4.3 Analysis of Anchors
@@
Table 3: Performance (%) of SRD module improves MVT [18], BUTD-DETR [21] and EDA [42] on ScanRefer [6] dataset.
@@ 4.5 Ablation Study
Analysis of ViewSRD Components. To assess the contribution of each component within ViewSRD, we conducted detailed ablation studies on the Nr3D dataset [1]. Starting from the full model, we systematically removed key modules one at a time to evaluate their individual impact. The results, presented in Table 4, demonstrate that each component plays a crucial role in enhancing model performance across different scenarios. Notably, the removal of the CCVTs leads to the most significant performance degradation. This is primarily because, without the view token, the model lacks explicit viewpoint information, impairing its ability to distinguish between different perspectives. Similarly, removing either the textual module or the scene module results in a noticeable decline, underscoring the necessity of cross-modal interaction. When view-alignment weighting is disabled (w/o Weight), performance drops by 0.9%, showing that dynamic alignment of view features is critical for performance under view-dependent conditions. Removing the SRD module leads to performance degradation, confirming the benefit of multi-anchor query decoupling. We also compare it with an LLM-based augmentation method from Multi3DRefer [52] and find that SRD achieves greater gains, highlighting the advantage of structured query decomposition over generic augmentation.
Analysis of Multi-View Modeling. We evaluate the effect of varying view counts on 3DVG performance using the Nr3D dataset. As shown in Table 5, testing with more views consistently improves accuracy when the model is trained with four views, highlighting the benefit of aggregating complementary spatial cues from multiple perspectives. When training and testing with the same number of views, performance improves from 64.4% (1 view) to 67.7% (2 views), but plateaus at 68.4% with 8 views, suggesting diminishing returns. Notably, four views offer a strong trade-off, capturing diverse spatial information with minimal redundancy and maintaining computational efficiency. This also suggests that uniformly attending to many views may dilute focus on key perspectives. Future work will explore adaptive view selection.
@@
5 Conclusion
A limitation of ViewSRD is its assumption that complex queries can be fully decomposed without overlapping relationships. While the decomposition into overlapping relations does not degrade performance, it diminishes the intended benefits of simplification. Future work will explore adaptive query to better preserve contextual dependencies.
Appendix
Appendix A Summary
This supplementary material offers detailed technical information on key components of our proposed framework, which were abbreviated in the main text. Specifically, we elaborate on the following: the target classifier $Clas$, the examples in template and the details of the sentence matching algorithm in Section 3.1; details of the loss function in Section 3.5.
Appendix B Target Classifier $Clas$
The target label classifier, denoted as $Clas$, is designed to regress the target’s class label from an input query $Q$. By determining the category of $Q$, we then determine which word in $Q$ belongs to that category to determine which word to use as the Target. Essentially, $Clas$ processes an input sentence by first encoding it via a pre-trained BERT [11] model to obtain a semantic embedding, and then maps this embedding through a multi-layer perceptron (MLP) to yield a probability distribution over $C$ target classes.
Specifically, given an input query $Q$, we first compute its embedding:
$F_{Q}=\mathrm{BERT}(Q),$ (8)
where $F_{Q}\in\mathbb{R}^{K\times d}$ represents the BERT output for the query $Q$, with $K$ being the number of tokens in $Q$ and $d$ denoting the feature dimension. We select the embedding corresponding to the [CLS] token as the aggregate representation $h$. This $h$ is then fed into the MLP-based classification head of $Clas$, implemented as a fully connected layer that produces logits $z\in\mathbb{R}^{C}$. These logits are subsequently converted into a probability distribution over the $C$ classes via the softmax function:
$P(c\mid Q)=\frac{\exp(z_{c})}{\sum_{j=1}^{C}\exp(z_{j})},$ (9)
where $z_{c}$ denotes the $c$-th element of $z$.
The classifier $Clas$ is trained by minimizing the standard cross-entropy loss:
$\mathcal{L}_{\text{cls}}=-\sum_{c=1}^{C}y_{c}\log\big{(}P(c\mid Q)\big{)},$ (10)
where $y_{c}$ is the one-hot encoded ground-truth label for class $c$. During inference, the predicted target label is given by:
$T_{lab}=\arg\max_{c}\,P(c\mid Q).$ (11)
In summary, $Clas$ effectively bridges the input query and the target object label by leveraging BERT-based encoding followed by MLP-based classification, ensuring robust target label regression in language-conditioned 3D grounding.
Appendix C Examples in SRD Module’s Template
In our Simple Relation Decoupling (SRD) module, different examples help LLMs in comprehending various descriptions of spatial relationships. To emphasize those relationships that are particularly sensitive to viewpoint changes, we select spatial relations whose interpretation can vary significantly with the observer’s perspective. All examples we used are as follows:
  • The {target} is close to the {anchor}.
  • The {target} is far away from the {anchor}.
  • The {target} is on the left of the {anchor}.
  • The {target} is on the right of the {anchor}.
  • The {target} is in front of the {anchor}.
  • The {target} is behind the {anchor}.
Appendix D Sentence Matching Algorithm
As mentioned in Section C, we predefine an Example Set $E$, which consists of $k$ examples in total. Given an anchor, it returns $k$ sentences. We employ a Sentence Matching Algorithm, as outlined in Algorithm 1, to select the sentence that best aligns semantically with the original sentence. Specifically, after the input of the target, a single anchor, and $k$ examples, we obtained $k$ candidate sentences, denoted as $S=\{{S_{1},S_{2},\ldots,S_{k}}\}$, and the original complex query is denoted as $S_{q}$ through LLMs. All sentences are then predicted through the Target Classifier $Clas$, to obtain pseudo-label $\hat{P}_{q}$ and $\hat{P}_{i},i\in\{1,2,\ldots,k\}$. The matching score consists of two components. The first is the consistency score, which is the evaluation score for maintaining consistency of the pseudo-label between the generated sentence and the query. The corresponding $S_{consist}$ is formulated as follows:
[…]
Algorithm 1 Sentence Matching Algorithm
1:  Input: Utterance $U$, Anchor set $O$, Target label $y$
2:  $N_{q}$ = len(U), $N_{anchor}$ = len(O)
3:  For each anchor in Anchor set $O$ do:
4:      For each example in Example set $E$ do:
[…]
6:       sentences = LLM(query)
[…]
8:       Predict labels = $Clas$(token)
9:       $S_{consist}$ = GetConsistScore(Predict labels)
10:       $S_{similar}$ = GetSimilarScore(sentences,U)
[…]
12:       Update the simplified sentence and Max_Score
13:       End for
14:      Anchor_LLMsentence.append(simplified sentence)
15:  End for
16:  Output: Sentence with Maximum Similarity
To preserve the semantic integrity of the generated sentence while preventing redundant generation from LLMs, we introduce the second part of the evaluation score: the similarity score. This score primarily considers the shared word count and the lengths of the sentences. We break down both sentences into individual words and calculate the number of shared words $N_{co\_word}$ between $S_{i}$ and $S_{q}$. To counteract the advantage of longer sentences in similarity calculation, we employed TextRank Algorithm[32] to compute the sentence similarity, denoted as $S_{Similar}$. Finally, the similarity score is calculated by dividing the shared word count by the sum of the logarithms of the lengths of the two sentences. The formula is as follows:
[…]
where $N_{q}$ and $N_{s}$ denote the length of the query and the generated sentence. In addition, we have defined an ideal decoupling sentence length $N^{*}$ based on the length of the sentence and the number of anchors in the sentence. Using $N^{*}$ as the center, the corresponding weights are obtained with a peak value 1, gradually decaying towards both ends. For instance, if $N_{q}$ is 7 and there are 2 anchors, the weight probability distribution would be {0.4, 0.6, 0.8, 1, 0.8, 0.6, 0.4}. When $N_{s}=2$, the corresponding weight probability $w$ is 0.6. Therefore, the corresponding formula for the score is as follows:
$\text{Score }=w\cdot S_{Similar}+S_{consist}.$ (14)
Appendix E Loss Function
Object-level Loss $\mathcal{L}_{\text{Object}}$: Followed by MVT[18], semantic class representations $\{L_{c}\}$ are obtained by encoding class name tokens with a pre-trained language encoder and extracting the [CLS] token. Given projected object features $F^{\prime}\in\mathbb{R}^{N\times D}$ (for $N$ objects with feature dimension $D$) and class representations $L\in\mathbb{R}^{C\times D}$, we compute the predicted logits as
$P_{\text{obj}}=F^{\prime}\cdot L^{T}\in\mathbb{R}^{N\times C}.$ (15)
The ground truth is represented as a one-hot encoded matrix $T_{\text{obj}}\in\mathbb{R}^{N\times C}$. The object-level loss is then defined as
$\mathcal{L}_{\text{Object}}=-\sum_{j=1}^{N}\sum_{c=1}^{C}T_{\text{obj}}(j,c)\,\log\Bigl{(}\text{softmax}(P_{\text{obj}})_{jc}\Bigr{)}.$ (16)
This loss ensures that the object features capture critical geometric properties, such as shape and center, and align accurately with their semantic categories.
Referential Loss $\mathcal{L}^{P}_{\text{ref}}$: To supervise spatial alignment, followed by CoT3DRef[2], , we obtain in parallel the anchor position labels $\mathbf{T}_{\text{anchor}}\in\mathbb{R}^{I}$ and the target position label $\mathbf{T}_{\text{target}}\in\mathbb{R}^{1}$, and concatenate them to form the referential ground truth
$\mathbf{T}_{\text{ref}}=\operatorname{concat}\left(\mathbf{T}_{\text{anchor}},\,\mathbf{T}_{\text{target}}\right)\in\mathbb{R}^{(I+1)},$ (17)
where each element $T_{\text{ref},p}$ (for $p=1,\ldots,I+1$) indicates the correct object index (among the $i$ objects present in the scene) for that spatial position. After view aggregation and prediction head, the model outputs referential logits $L_{\text{ref}}\in\mathbb{R}^{(I+1)\times i}$, where each row corresponds to a spatial position (anchor or target) and each column represents one of the $i$ objects. The referential loss is defined as
$\mathcal{L}^{P}_{\text{ref}}=-\frac{1}{I+1}\sum_{p=1}^{I+1}\log\frac{\exp\big{(}L_{\text{ref},p,T_{\text{ref},p}\big{)}}}{\sum_{r=1}^{i}\exp\big{(}L_{\text{ref},p,r}\big{)}}.$ (18)
This loss encourages the textual module to generate discriminative features that accurately capture the semantic nuances of the grounding statement, effectively distinguishing between the anchor and target labels.
Sentence-level Loss $\mathcal{L}_{\text{Sent}}$: Followed by CoT3DRef[2], we supervise the textual module using a standard cross-entropy loss. Let $I$ denote the number of anchor class labels in the original sentence. For each sample, we form a label sequence by concatenating the $I$ anchor labels with one target label, yielding a sequence of length
$S=I+1.$ (19)
After Textual Aggregation, the classification head produces language prediction logits $\mathbf{P}_{\text{sent}}\in\mathbb{R}^{S\times C}$, where $C$ is the total number of classes. The ground truth is represented as a vector $\mathbf{T}_{\text{sent}}\in\mathbb{R}^{S}$, with each element $T_{\text{sent},i}$ indicating the correct class index for the $i$-th position. The language loss is defined as
$\mathcal{L}_{\text{Sent}}=-\frac{1}{S}\sum_{i=1}^{S}\log\frac{\exp\big{(}(\mathbf{P}_{\text{sent}})_{i,T_{\text{sent},i}}\big{)}}{\sum_{j=1}^{C}\exp\big{(}(\mathbf{P}_{\text{sent}})_{ij}\big{)}}.$ (20)
This loss encourages the textual module to generate discriminative features that accurately capture the semantic nuances necessary for language grounding.
Appendix F t-SNE visualization of CCVTS
Figure 5: t-sne Visualization of CCVTs.
To further investigate how CCVTs encode perspective information, we visualize the learned features using t-SNE, as shown in Fig. 5. Tokens originating from different viewpoints form well-separated clusters, while those from the same viewpoint are closely grouped. This clear structural separation illustrates that CCVTs effectively capture and preserve view-specific semantics, which are essential for accurate cross-view alignment in 3D visual grounding.
Appendix G Analysis of Loss Weight.